A loan of $5000 must be repaid in 10 years at 4.75% compounded quarterly.
A) How much will have to be paid at the end of the 10 years?
B) How much interest will have to be paid?

Answer by Theo(13342)

the loan is 5000.
it must be repaid in 10 years at 4.75% compounded quarterly.
4.75% per year compounded quarterly = 4.75/4 = 1.1875% per quarter.
number of quarters of a year in 10 years = 40.
formula to use is f = p * (1 + r) ^ n
f is the future value
p is the present value
n is the number of time periods
r is interest rate per time period.
the formula uses the interest rate, not the percent.
1.1875% = .011875.
the formula becomes:
f = 5000 * (1 + .011875) ^ 40
solve for f to get f = 8017.604852.
that's how much will have to be paid at the end of the 10 years.
the interest portion of that payment would be that minus 5000 = 3017.604852.
let me know if you have any questions.
theo